Fill in the blanks in the following separate income statements a through e. (Amounts to be deducted should be indicated by a minus sign.) 

$ 64,000 $ 44,800 $ 39,000 $ 24,200 Cost of goods sold 16,720 4.480 7,300 32,000 5,800 . (8,900) 7,300 33,000 (7.600) 6.700 Merchandise inventory (beginning) Total cost of merchandise purchases Merchandise inventory (ending) Cost of goods sold Gross profit Expenses Net income (loss) (2,800) 17.000 33,630 17.000 | $ 4,300 11,820 (7.520) 11.890 15,910 6.700 48.700 5.700 43,000 $ $
Answer #1

ANSWER:

Cost of goods sold=Beginning inventory+Purchases-Ending inventory
Sales-COGS=Gross profit
Gross profit-Expenses=Net income
Calculation a b c d e
Sales A      64,000       44,800       39,000      81,400      24,200
COGS:
Beginning inventory B        7,300       16,720         5,800        7,300        4,480
Purchases C      32,000         3,080       37,800      33,000        6,700
Ending inventory D      -5,670        -2,800       -8,900      -7,600       -3,480
COGS F=B+C-D      33,630       17,000       34,700      32,700        7,700
Gross profit G=A-F      30,370       27,800         4,300      48,700      16,500
Expenses H      17,000       11,890       11,820        5,700        6,700
Net income I=G-H      13,370       15,910       -7,520      43,000        9,800